# Statistical evaluation and optimization of entanglement purification protocols

### v1

@@ V Summary and conclusions
To summarize, we have presented a numerical method, that is capable of characterizing the performance of entanglement purification protocols. We have presented a new CNOT protocol, which was evaluated together with two pioneering protocols and a recent proposal based on matter-field interactions. Our results show that all the protocols destroy entanglement in the very first iteration. This was known for the pioneering protocols, and here in addition we have demonstrated quantitatively that only a small set of entangled states are kept. Even though these states are turned into a Bell state in the limit of infinite iterations, the average concurrence of the whole sample stays below its starting value. The MFI-based and the new CNOT protocols perform better and they can turn slightly more than $21\%$ of the two-qubit states into a Bell state.
@@
We have defined the ultimate limit of all possible purification protocols, which is nothing else than the percentage of all entangled states within the set of all quantum states, i.e., approximately $75\%$. In other words, an ultimate protocol can purify all entangled states into a maximally entangled one. In this context, we have searched for optimal asymmetric entangling gates in the nodes $A$ and $B$. We have found that this approach is improving the average concurrence very slowly. Therefore, motivated by the other approaches, we have included in the first iteration an entangling projection, which destroys some entanglement. This strategy turns out to be a boost for the optimized variational approach, which can outperform all protocols discussed in this manuscript. However, even the variational approach seems unable to reach the ultimate bound for entanglement purification. At the current stage, we cannot determine whether this result could be improved by implementing different quantum operations or optimization methods. There might also be an upper bound for this family of protocols that lies below the ultimate asymptote.
@@
Our numerical analysis focuses on the improvement of the average concurrence. However, we would like point out that all the protocols investigated in this work have a common property. If some information about the input state is known, e.g., the state has an overlap strictly larger than $0.5$ with one of the Bell states, then these protocols convert purifiable entangled states into a fixed maximally entangled state, which is usually a Bell state.
@@
Finally, some comments on the average success probabilities are in order. It is known that improving the concurrence alone is not good enough measure, because success probabilities play a crucial role in the identification of required resources, i.e., how many qubits are required to perform some iterations. To improve the success probabilities as well, this leads to a multiobjective optimization task. This is not included here, because this work aims to introduce a general evaluation scheme based on the hit-and-run algorithm and the concurrence, which have given an insight into the performance of entanglement purification protocols and showed that improvements are possible in computer-based protocol designs.
@@ Acknowledgements.
This work was supported by AIDAS-AI, Data Analytics and Scalable Simulation, which is a Joint Virtual Laboratory gathering the Forschungszentrum Jülich and the French Alternative Energies and Atomic Energy Commission, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534, and by the Jülich Supercomputing Center. We are grateful to J. M. Torres and T. Calarco for stimulating discussions. Simulations were realized in PYTHON using the library JAX [41] and in JULIA. The code and the data are available at https://github.com/franz3105/soepp.
@@ Appendix A Hit-and-run algorithm
In this Appendix, we briefly describe the steps of the hit-and-run algorithm in Algorithm 2. Given a $K\subseteq\mathbb{R}^{15}$, we generate for an ${\bf a}\in K$ a random uniform vector ${\bf x}$ on the sphere, which is around ${\bf a}$ and has unit radius. We generate a random uniform number $\lambda$ on the interval $[-\sqrt{3}/2,\sqrt{3}/2]$, because $K$ is inside the sphere of radius $\sqrt{3}/2$ around the origin [17]. If ${\bf a}^{\prime}={\bf a}+\lambda{\bf x}\in K$, then we move there, otherwise we start all over from ${\bf a}$. We always start the sampling from ${\bf a}={\bf 0}$, i.e., the maximally mixed state.
@@
Figure 3: Unbiased sample variances of the concurrence and the success probability as a function of the number of iterations for the non-variational (a), (b) and the optimized variational protocols (c), (d). We use the same color scheme as in Figs. 1 and 2. We observe that the sample variance grows with the number of iterations and reaches its maximum asymptotic values when all the density matrices are all mapped to either one or zero concurrence (see Eq. (62)). The fraction of density matrices with non-zero concurrence in the limit of infinite iterations determines the performance of the protocol and also the maximum of the sample variance.
@@ Appendix B Statistics of entanglement purification
Figure 4: (a) Bar chart of the distribution of the concurrence for a sample of two-qubit density matrices after removing the states with concurrence zero; and (b) average initial concurrence as a function of the number of samples generated by the hit-and-run Monte Carlo algorithm. Vertical bars show the standard error of the mean. We see that the average approximately stabilizes after $N=10^{3}$ samples. This is only true for the initial concurrence, as the average concurrence after each run of the purification protocol need a higher resolution (empirically we find $N>10^{4}$, especially for the MFI protocol). The first plot (a) uses $10^{6}$ density matrices, obtained by running the hit-and-run algorithm once.
In this Appendix, details concerning the sample mean and standard deviation are discussed. The primary task is to estimate numerically different averages of the concurrence or the success probability over all two-qubit density matrices. In the main text, we have already identified this set with the convex body $K$ in the Euclidean space $\mathbb{R}^{15}$. Thus, every density matrix $\rho$ in Eq. (38) is uniquely described by a vector ${\bf a}\in K$. In this context, the average concurrence reads
@@
i.e., the average concurrence is always smaller than $S/N$ (see Eq. (61)). Then the sample standard deviation reads
@@
$\epsilon_{i}=\sum^{S}_{j=1}\epsilon^{(i)}_{j}\quad\text{and}\quad\beta_{i}=\sum^{S}_{j=1}\left(\epsilon^{(i)}_{j}\right)^{2}.$ (66)
@@
Figure 5: Bar charts of the concurrence distribution for the Bennett – (a), (b), (c) –, Deutsch – (d), (e), (f), and optimized protocol (using the argmax policy in Eq. (34) and the projector $\Pi$ in Eq. (58)) – (g), (h), (i) –, each one for the second, seventh, and fifteenth iteration – corresponding to the first, second and third column of plots. States with concurrence zero have been removed to allow us to visualize the action of the protocols, as they skew the distribution due to their generally high number. Vertical axes have been set so to range between 0 and $6\times 10^{4}$, so that the reader can visualize the growth of the number of states with concurrence equal to one on the right as the number of iterations grows. We see that the optimized protocol preserves the most states, whereas the number of purified states for Deutsch and Bennett is significantly lower. These bar charts use $N=10^{6}$ samples of the concurrence (one run of the hit-and-run algorithm).
In Fig. 5, we see the iteration-based evolution of the distribution of the concurrence for $\mathcal{C}>0$. Each column represents one of three different points in the purification protocols: (a), (d), (g) $i=1$; (b), (e), (h) $i=7$ and (c), (f), (i) $i=14$. We show here nine plots, where the first row (a), (b), (c) represents the Bennett protocol, the second row (d), (e), (f) the MFI protocol and the third one (g), (h), (i) the CNOT protocol. We find, as expected, that the number of states with non-zero concurrence is much lower in the case of the Bennett protocol than for the other two, although every protocol improves the distribution of the concurrence towards $\bar{\mathcal{C}}=1$.
@@ Appendix C Statistical evaluation of the fidelities
where the states $k=1,2,3,4$ are the Bell states defined in Eq. (II). We then use them to define the output fidelity for the Bennett protocol [2]
@@
[…]
@@
[…]
Output fidelities need to be computed according to Eqs. (70), (71), (72), (73), (74), (75) and (76) for each sampled density matrix at each iteration of the corresponding protocol. The average fidelity is defined as
@@
The average fidelities for the purifiable states with respect to all stable fixed points for the Bennett, Deutsch, MFI, and CNOT protocols are given in Fig. 6. We see that the number of states that can be brought to a stable fixed point by the MFI and CNOT protocols is significantly larger than those of Bennett and Deutsch. This is in accordance with the findings presented in the main text.
Prior knowledge of the input state is essential because most of the protocols currently available in the literature can only work if one knows that the states to be purified have certain underlying properties, which need to be known to avoid failure. In the subsequent discussion, we are going to elucidate this argument by a simple demonstration. Let us now consider a general two-qubit density matrix that we wish to purify towards a Bell state. The Bennett protocol starts with a twirling operation, whose only goal is to bring the state $\rho$ into the Werner form
@@
This operation, however, does not guarantee the success of the purification because the Bennett protocol can only work if the condition $r_{1}>0.5$ is fulfilled. If it does, then the Bennett protocol in the asymptotic limit brings this state to the Bell state |$\ket{1}$. In all other cases, the purification protocol fails and leads to a mixed state. The infinite-limit output of the Bennett protocol for a general two-qubit density matrix is therefore a classical mixture of the purified states (with concurrence one) and the mixed states for which the protocol failed (with concurrence zero). It turns out that for most of the density matrices, the protocol fails, i.e. the twirling operation maps them to Werner states with $r_{1}<0.5$, which is why the blue curve in Fig. 1 has so small values. The asymptotic value of the curve represents exactly the fraction of states that can be purified. The output of the Bennett protocol with asymptote $\bar{\mathcal{C}}^{(\infty)}_{\text{Bennett}}$ for a general random density matrix in the limit of infinite iterations will be approximately
$\rho_{\text{Bennett}}=\bar{\mathcal{C}}^{(\infty)}_{\text{Bennett}}\ket{1}\bra{1}+\frac{(1-\bar{\mathcal{C}}^{(\infty)}_{\text{Bennett}})}{3}(I_{4}-\ket{1}\ket{1})$ (79)
@@
Figure 6: Average output fidelities (calculated using Eqs. (70), (71), (72), (73), (74), (75), (76), and (77) for the different fixed points) with respect to the attractors of the different protocols as a function of the number of iterations. For the Bennett protocol, |$\ket{1}$ is the only fixed point. For the other protocols, the attractors are, respectively, |$\ket{2}$ and |$\ket{4}$ (Deutsch), |$\ket{1}$ and |$\ket{2}$ (MFI) and |$\ket{4}$ and |$\ket{2}$ (CNOT). We see that for each of their attractors, the share of purifiable states of both the MFI and the CNOT protocols are significantly larger than the ones of Bennett and Deutsch. The plot is obtained by running $10$ simulations of the hit-and-run algorithm, each one with $N=10^{6}$ samples.
The Deutsch protocol, unlike Bennett, has three stable fixed points with two Bell states, which is often not mentioned in the literature [1], but after the appearance of the proposal this has been thoroughly investigated in Ref. [29]. Now, one obtains a classical mixture of three states with weights defined by $\bar{\mathcal{C}}^{(\infty)}_{\text{Deutsch}}$, where the states |$\ket{2}\bra{2}$ | and |$\ket{4}\bra{4}$ | have the same weights as it is shown Fig.  1. This leads also to a noisy entangled state. In conclusion, one cannot start with an unknown density matrix and just run the protocols without having some prior information beforehand, because, as we see, the protocols mostly fail if certain conditions are not met. These conditions, as we have shown, may be very restrictive or more relaxed, in which case a broader class of input states can be successfully purified.

### v2

@@ V Summary and conclusions
To summarize, we have presented a numerical method that is capable of characterizing the performance of entanglement purification protocols. We have presented a CNOT-based protocol, which was evaluated together with two pioneering protocols and a recent proposal based on matter-field interactions. Our results show that all the protocols destroy entanglement in the very first iteration. This was known for the pioneering protocols, and here in addition we have demonstrated quantitatively that only a small set of entangled states is kept. Even though these states are turned into a Bell state in the limit of infinite iterations, the average concurrence of the whole sample stays below its starting value. The MFI-based and the CNOT-based protocols perform better and they can turn slightly more than $21\%$ of the two-qubit states into a Bell state.
@@
We have defined the ultimate limit of all possible purification protocols, which is nothing other than the percentage of all entangled states within the set of all quantum states, i.e., approximately $75\%$. In other words, an ultimate protocol can purify all entangled states into maximally entangled ones. In this context, we have searched for optimal asymmetric entangling gates in the nodes $A$ and $B$. We have found that this approach is improving the average concurrence very slowly. Therefore, motivated by the other approaches, we have included in the first iteration an entangling projection, which destroys some entanglement. This strategy turns out to be a boost for the optimized variational approach, which can outperform all protocols discussed in this paper. However, even the variational approach seems unable to reach the ultimate bound for entanglement purification. At the current stage, we cannot determine whether this result could be improved by implementing different quantum operations or optimization methods. There might also be an upper bound for this family of protocols that lies below the ultimate asymptote.
@@
Our numerical analysis focuses on the improvement of the average concurrence. However, we would like to point out that all the protocols investigated in this work have a common property. If some information about the input state is known, e.g., the state has an overlap strictly larger than $0.5$ with one of the Bell states, then we know beforehand whether or not these protocols convert entangled states into a fixed maximally entangled state, which is usually a Bell state. A further difficulty is that the iterations of the protocol are nonlinear quantum state transformations, which lead to chaotic behavior [41, 42, 43, 44]. Therefore, in the trace norm topology [38], seemingly close density matrices might have different future trajectories. Our approach avoids this interesting but complicated behavior by using the average concurrence, a choice, which we have demonstrated to be also successful in finding different and effective entanglement purification protocols.
@@
Finally, some comments on the average success probabilities are in order. It is known that improving the concurrence alone is not a good enough measure, because success probabilities play a crucial role in the identification of required resources, i.e., how many qubits are required to perform some iterations. To improve the success probabilities as well, this leads to a multiobjective optimization task. This is not included here, because this work aims to introduce a general evaluation scheme based on the hit-and-run algorithm and the concurrence, which have given insight into the performance of entanglement purification protocols and shown that improvements are possible in computer-based protocol designs.
@@ Acknowledgements.
This work was supported by AIDAS-AI, Data Analytics and Scalable Simulation, which is a Joint Virtual Laboratory gathering the Forschungszentrum Jülich and the French Alternative Energies and Atomic Energy Commission, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 – 390534, and by the Jülich Supercomputing Center. We are grateful to J. M. Torres and T. Calarco for stimulating discussions. Simulations were realized in PYTHON using the library JAX [26] and in JULIA. The code and the data are available at https://github.com/franz3105/soepp (code) and https://zenodo.org/uploads/10996551 (data sets).
@@ Appendix A Hit-and-run algorithm
In this Appendix we briefly describe the steps of the hit-and-run algorithm in Algorithm 2. Given a $K\subseteq\mathbb{R}^{15}$, we generate for an ${\bf a}\in K$ a random uniform vector ${\bf x}$ on the sphere, which is around ${\bf a}$ and has unit radius. We generate a random uniform number $\lambda$ on the interval $[-\sqrt{3}/2,\sqrt{3}/2]$, because $K$ is inside the sphere of radius $\sqrt{3}/2$ around the origin [17]. If ${\bf a}^{\prime}={\bf a}+\lambda{\bf x}\in K$, then we move there, otherwise we start all over from ${\bf a}$. We always start the sampling from ${\bf a}={\bf 0}$, i.e., the maximally mixed state.
@@
Figure 3: Unbiased sample variances of the concurrence and the success probability as a function of the number of iterations for (a) and (b) the non-variational and (c) and (d) the optimized variational protocols. We use the same color scheme as in Figs. 1 and 2. We observe that the sample variance grows with the number of iterations and reaches its maximum asymptotic values when all the density matrices are all mapped to either one or zero concurrence [see Eq. (65)]. The fraction of density matrices with non-zero concurrence in the limit of infinite iterations determines the performance of the protocol and also the maximum of the sample variance.
@@ Appendix B Statistics of entanglement purification
Figure 4: (a) Bar chart of the distribution of the concurrence for a sample of two-qubit density matrices after removing the states with concurrence zero; and (b) average initial concurrence as a function of the number of samples generated by the hit-and-run Monte Carlo algorithm. Vertical bars show the standard error of the mean. We see that the average approximately stabilizes after $N=10^{3}$ samples. This is only true for the initial concurrence, as the average concurrence after each run of the purification protocol needs a higher resolution (empirically we find $N>10^{4}$, especially for the MFI protocol). Plot (a) uses $10^{6}$ density matrices, obtained by running the hit-and-run algorithm once.
In this Appendix, details concerning the sample mean and standard deviation are discussed. The primary task is to estimate numerically different averages of the concurrence or the success probability over all two-qubit density matrices. In the main text we have already identified this set with the convex body $K$ in the Euclidean space $\mathbb{R}^{15}$. Thus, every density matrix $\rho$ in Eq. (41) is uniquely described by a vector ${\bf a}\in K$. In this context, the average concurrence reads
@@
i.e., the average concurrence is always smaller than $S/N$ [see Eq. (64)]. Then the sample standard deviation reads
@@
$\epsilon_{i}=\sum^{S}_{j=1}\epsilon^{(i)}_{j},\quad\beta_{i}=\sum^{S}_{j=1}\left(\epsilon^{(i)}_{j}\right)^{2}.$ (69)
@@
Figure 5: Bar charts representing 50 bins of the concurrence distribution for the Bennett – (a), (b), (c), Deutsch – (d), (e), (f), and optimized protocol [using the argmax policy in Eq. (39) and the projector $\Pi$ in Eq. (61)] – (g), (h), (i), each one for the second, seventh, and fifteenth iteration – corresponding to the first, second and third column of the plots. States with concurrence zero have been removed to allow us to visualize the action of the protocols, as they skew the distribution due to their generally high number. As a consequence, the bar charts exhibit different heights, since the three protocols map different numbers of states to concurrence zero. Vertical axes have been set so to range between 0 and $6\times 10^{4}$, so that the reader can visualize the growth of the number of states with concurrence equal to one on the right as the number of iterations grows. We see that the optimized protocol preserves the most states, whereas the number of purified states for the Deutsch and Bennett protocols is significantly lower. These bar charts use $N=10^{6}$ samples of the concurrence (one run of the hit-and-run algorithm).
In Fig. 5, we see the iteration-based evolution of the distribution of the concurrence for $\mathcal{C}>0$. Each column represents one of three different points in the purification protocols: (a), (d), (g) $i=2$; (b), (e), (h) $i=7$ and (c), (f), (i) $i=15$. We show here nine plots, where the first row (a), (b), (c) represents the Bennett protocol, the second row (d), (e), (f) the MFI protocol and the third one (g), (h), (i) the CNOT protocol. We find, as expected, that the number of states with non-zero concurrence is much lower in the case of the Bennett protocol than for the other two, although every protocol improves the distribution of the concurrence towards $\bar{\mathcal{C}}=1$.
@@ Appendix C Statistical evaluation of the fidelities
where the states $k=1,2,3,4$ are the Bell states defined before Eq. (II). We then use them to define the output fidelity for the Bennett protocol [2]
@@
$F^{(1)}_{\text{MFI}}(\rho)=\begin{cases}r_{1}&\text{\footnotesize if $(2r_{1}-1)(1-2r_{3})>-(2{\rm Im[r_{13}]})^{2}-(2{\rm Re}[r_{24}])^{2}$},\\ 0&\text{otherwise}\end{cases}$ (76)
$F^{(2)}_{\text{MFI}}(\rho)=\begin{cases}r_{2}&\text{\footnotesize if $(2r_{2}-1)(1-2r_{4})>-(2{\rm Im[r_{24}]})^{2}-(2{\rm Re}[r_{13}])^{2}$},\\ 0&\text{otherwise},\end{cases}$ (77)
@@
$F^{(4)}_{\text{CNOT}}(\rho)=\begin{cases}r_{4}&\text{\footnotesize if $(2r_{1}-1)(1-2r_{4})>-(2{\rm Im[r_{23}]})^{2}-(2{\rm Re}[r_{14}])^{2}$}\\ 0&\text{otherwise}\end{cases}$ (78)
$F^{(2)}_{\text{CNOT}}(\rho)=\begin{cases}r_{2}&\text{\footnotesize if $(2r_{2}-1)(1-2r_{3})>-(2{\rm Im[r_{14}]})^{2}-(2{\rm Re}[r_{23}])^{2}$}\\ 0&\text{otherwise}.\end{cases}$ (79)
Output fidelities need to be computed according to Eqs. (73)–(79) for each sampled density matrix at each iteration of the corresponding protocol. The average fidelity is defined as
@@
The average fidelities for the purifiable states with respect to all stable fixed points for the Bennett, Deutsch, MFI, and CNOT protocols are given in Fig. 6. We see that the number of states that can be brought to a stable fixed point by the MFI and CNOT protocols is significantly larger than those of the Bennett and Deutsch protocols. This is in accordance with the findings presented in the main text.
Prior knowledge of the input state is essential because most of the protocols currently available in the literature can only work if one knows that the states to be purified have certain underlying properties, which need to be known to avoid failure. In the subsequent discussion, we elucidate this argument by a simple demonstration. Let us now consider a general two-qubit density matrix that we wish to purify towards a Bell state. The Bennett protocol starts with a twirling operation, whose only goal is to bring the state $\rho$ into the Werner form
@@
This operation, however, does not guarantee the success of the purification because the Bennett protocol can only work if the condition $r_{1}>0.5$ is fulfilled. If it does, then the Bennett protocol in the asymptotic limit brings this state to the Bell state |$\ket{1}$. In all other cases, the purification protocol fails and leads to a mixed state. The infinite-limit output of the Bennett protocol for a general two-qubit density matrix is therefore a classical mixture of the purified states (with concurrence one) and the mixed states for which the protocol failed (with concurrence zero). It turns out that for most of the density matrices, the protocol fails, i.e. the twirling operation maps them to Werner states with $r_{1}<0.5$, which is why the blue curve in Fig. 1 has such small values. The asymptotic value of the curve represents exactly the fraction of states that can be purified. The output of the Bennett protocol with asymptote $\bar{\mathcal{C}}^{(\infty)}_{\text{Bennett}}$ for a general random density matrix in the limit of infinite iterations will be approximately
$\rho_{\text{Bennett}}=\bar{\mathcal{C}}^{(\infty)}_{\text{Bennett}}\ket{1}\bra{1}+\frac{(1-\bar{\mathcal{C}}^{(\infty)}_{\text{Bennett}})}{3}(I_{4}-\ket{1}\bra{1})$ (82)
@@
Figure 6: Average output fidelities [calculated using Eqs. (73)–(79), and (80) for the different fixed points] with respect to the attractors of the different protocols as a function of the number of iterations. For the Bennett protocol, |$\ket{1}$ is the only fixed point. For the other protocols, the attractors are |$\ket{2}$ and |$\ket{4}$ (Deutsch), |$\ket{1}$ and |$\ket{2}$ (MFI) and |$\ket{4}$ and |$\ket{2}$ (CNOT). We see that for each of their attractors, the shares of purifiable states of both the MFI and the CNOT protocols are significantly larger than the ones of Bennett and Deutsch. The plot is obtained by running ten simulations of the hit-and-run algorithm, each one with $N=10^{6}$ samples.
The Deutsch protocol, unlike the Bennett, has three stable fixed points with two Bell states, which is often not mentioned in the literature [1], but after the appearance of the proposal this has been thoroughly investigated in Ref. [28]. Now, one obtains a classical mixture of three states with weights defined by $\bar{\mathcal{C}}^{(\infty)}_{\text{Deutsch}}$, where the states |$\ket{2}\bra{2}$ | and |$\ket{4}\bra{4}$ | have the same weights as it is shown in Fig. 1. This leads also to a noisy entangled state. In conclusion, one cannot start with an unknown density matrix and just run the protocols without having some prior information beforehand, because, as we see, the protocols mostly fail if certain conditions are not met. These conditions, as we have shown, may be very restrictive or more relaxed, in which case a broader class of input states can be successfully purified.